### v2

# Eka-Eval : A Comprehensive Framework for Indian Large Language Models

Abstract
The rapid advancement of Large Language Models (LLMs) has intensified the need for evaluation frameworks that address the requirements of linguistically diverse regions, such as India, and go beyond English-centric benchmarks. We introduce Eka-Eval, a unified evaluation framework that integrates over 35+ benchmarks (including 10 Indic benchmarks) across nine major evaluation categories. The Framework provides broader coverage than existing Indian language evaluation tools, offering 11 core capabilities through a modular architecture, seamless integration with Hugging Face and proprietary models, and plug-and-play usability. As the first end-to-end suite for scalable, multilingual LLM benchmarking, the framework combines extensive benchmarks, modular workflows, and dedicated support for low-resource Indian languages to enable inclusive assessment of LLM capabilities across diverse domains. We conducted extensive comparisons against five existing baselines, demonstrating that Eka-Eval achieves the highest participant ratings the highest participant ratings in four out of five categories.
[…]
Eka-Eval : A Comprehensive Framework for Indian Large Language Models
[…]
1 Introduction
Large Language Models (LLMs) have rapidly transformed natural language processing (NLP), enabling impressive generalization across diverse tasks including instruction following, reasoning, summarization, translation, and tool use. With the advent of general-purpose foundation models such as GPT-4 (Achiam et al., 2023), Claude (Anthropic, 2024), Gemini (Anil et al., 2023) and Llama-3 (Touvron et al., 2023), the focus of research has increasingly shifted from building task-specific models to systematically evaluating these powerful systems. Evaluation plays a critical role not only in measuring progress but in identifying capabilities, exposing limitations, and informing deployment strategies.
Figure 1: Architecture of Eka-Eval: A modular framework combining model inference, benchmark registry, evaluation engine, and result processing with support for caching and distributed inference.
In response, several evaluation frameworks have emerged, including HELM (Liang et al., 2022), EleutherAI Harness (Gao et al., 2021), lm-eval-harness (Gao et al., 2021), and OpenCompass (OpenCompass Contributors, 2023). However, existing evaluation frameworks are primarily centered on English or other high-resource languages and offer limited support for low-resource or multilingual settings (Watts et al., 2024). This lack of inclusivity significantly limits the effectiveness of these tools in linguistically diverse regions like India, which has 22 constitutionally recognized languages encompassing over a billion native speakers.
While notable benchmarks like IndicGLUE (Kakwani et al., 2020), IndicGenBench (Singh et al., 2024), and MILU (Verma et al., 2024) have addressed some gaps in Indic-language evaluation, they remain siloed efforts and lack integration into unified evaluation workflows — requiring users to perform manual configuration to use them in real-world LLM assessment pipelines.
Furthermore, popular frameworks such as FreeEval (Yu et al., 2024) and lm-eval-harness (Gao et al., 2021) require extensive configuration and engineering expertise, limiting their adoption by developers and researchers operating in low-resource environments. These challenges create a need for an actively maintained, community driven, multilingual, and task-diverse evaluation suite.
Framework Custom Datasets Custom Models Custom Prompting Long Context Tool Use Distributed Inference Visual Analysis Multilingual Indic Production Optimization Interactive CLI Quantization
lm-eval-harness
OpenCompass
HELM
OpenAI Evals
DeepEval
FreeEval
indic-eval
Eka-Eval
Table 1: Comparison of LLM Evaluation Frameworks across eleven key capabilities, showing the comprehensive coverage of Eka-Eval.
To address these limitations, we introduce Eka-Eval, a unified, extensible, and ready-to-use evaluation framework for LLMs, integrating over 35 benchmarks spanning both global (English) and Indic-language tasks. The overall architecture, presented in Figure 1, is designed to be modular, compatible with Hugging Face and proprietary models, it offers plug-and-play usability with minimal configuration overhead. To contextualize Eka-Eval’s capabilities, Table 1 compares leading LLM evaluation frameworks across eleven key capabilities, demonstrating its more comprehensive feature coverage relative to existing frameworks.
Eka-Eval covers nine major evaluation categories: (i) Code Generation and programming; (ii) Mathematics and logical reasoning; (iii) Reading comprehension; (iv) Commonsense reasoning; (v) World knowledge; (vi) Long-context understanding; (vii) General reasoning and knowledge; (viii) Tool use and API reasoning; and (viiii) Indic-specific NLP benchmarks including ARC-C (Clark et al., 2018), BoolQ (Clark et al., 2019), GSM8K (Cobbe et al., 2021), Math (Hendrycks et al., 2021), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2020a), SQuAD (Rajpurkar et al., 2018), APIBench (Patil et al., 2024), IndicGenBench (Singh et al., 2024).
Our key contributions are:
  • We propose Eka-Eval, a unified and modular evaluation framework that integrates over 35 benchmarks across nine major evaluation categories, providing comprehensive assessment of both global and Indic-language LLM capabilities.
  • We provide comprehensive support for most Indian languages and conduct thorough human evaluation across five existing frameworks on four benchmark tasks, demonstrating Eka-Eval’s effectiveness and reliability.
2 Related Work
The evaluation of LLMs has evolved from task-specific benchmarks to comprehensive, modular frameworks that support diverse capabilities and deployment settings. This section categorizes prior work into three major areas: general-purpose evaluation frameworks, specialized capability benchmarks, and multilingual/Indic evaluations. We conclude by highlighting how Eka-Eval integrates and advances these directions.
General-Purpose LLM Evaluation Frameworks: Early benchmark suites such as GLUE (Wang et al., 2018) and SuperGLUE (Wang et al., 2019) laid the groundwork for multi-task evaluation, but were primarily English-centric and biased toward high-resource languages. Frameworks like HELM  (Liang et al., 2022) expanded evaluation to multidimensional axes—accuracy, calibration, robustness, fairness, and efficiency—applied uniformly across 30+ models. Similarly, BIG-Bench (Srivastava et al., 2022) curated over 200 tasks via crowdsourced contributions, highlighting model generality but with limited support for multilingualism or real-world deployment constraints. In parallel, tooling infrastructure also emerged. lm-eval-harness (Gao et al., 2021) provided reproducible few-shot evaluation pipelines, while OpenAI’s Evals (OpenAI Contributors, 2023) offered YAML-based structured evaluation for internal alignment research.
Recent frameworks emphasize modularity, extensibility, and accessibility. For instance, OpenCompass OpenCompass Contributors (2023) integrates distributed inference and real-time leaderboards, though it lacks Indic language support. FreeEval Yu et al. (2024) introduces meta-evaluation and contamination detection, but lacks robust multilingual and Indic language support. DeepEval DeepEval Contributors (2024) supports long-context and tool-use tasks but lacks flexible multilingual and low-resource customization, and its default metrics are often too rigid.
Specialized Capability Benchmarks: To evaluate emerging LLM capabilities, task-specific frameworks and datasets have been proposed. Tool-use and agentic behavior are assessed in ToolBench (Qin et al., 2023) and API-Bank (Li et al., 2023) respectively, which evaluate model interactions with real-world APIs.
Long-context reasoning is explored in InfiniteBench (Zhang et al., 2024) and RULER (Hsieh et al., 2024), which evaluate memory and coherence over 100K+ token sequences. Mathematical reasoning and code generation are evaluated using datasets like GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021), respectively. While powerful, these benchmarks have limited scope and lack integration into unified evaluation pipelines.
Multilingual and Indic Language Evaluation: Multilingual evaluation has gained prominence through benchmark suites such as XTREME Hu et al. (2020), xP3 Muennighoff et al. (2022), and MEGA Ahuja et al. (2023), covering 40–70 languages. However, these benchmarks often rely on automatic translation, which may not reflect natural language usage or cultural nuances. FLORES Goyal et al. (2022) improved translation evaluation with high-quality parallel corpora for low-resource languages.
For Indian languages, IndicGLUE (Kakwani et al., 2020) and IndicXTREME (Kakwani et al., 2020) pioneered NLU benchmarking across major Indic languages. IndicGenBench (Singh et al., 2024) broadened the scope to generation tasks over 29 Indic languages, but remains a dataset collection rather than a complete evaluation framework, lacking system-level extensibility or plug-and-play usability. Frameworks such as indic-eval (IndicEval Contributors, 2024) wrap existing tools, like LightEval (LightEval, 2024), to support select Indic tasks. However, they offer limited extensibility and lack support for long-context tasks, tool-use evaluation, and custom prompting—features increasingly essential in real-world use cases (Chen et al., 2023).
Eka-Eval is a unified, extensible evaluation framework that integrates the breadth of multilingual and capability-centric benchmarks within a single, production-ready system. It supports a wide spectrum of evaluation settings, including both global and Indic benchmarks spanning 29+ Indian languages, designed for practical deployment and large-scale experimentation.
3 Capabilities of LLM Evaluation Frameworks
We identify eleven foundational capabilities that are critical for the design and continuous advancement of modern LLM evaluation frameworks. Several of these capabilities are described in earlier evaluation frameworks such as lm-eval-harness (Gao et al., 2021), OpenCompass (OpenCompass Contributors, 2023), FreeEval (Yu et al., 2024), and DeepEval (DeepEval Contributors, 2024). These eleven capabilities are:
(1) Custom Datasets:Support for loading and evaluating user-defined datasets beyond standard benchmark corpora.
(2) Custom Models: Compatibility with a wide range of models, including local checkpoints and API-hosted endpoints.
(3) Custom Prompting: Flexible, template-based prompting paradigms encompassing zero-shot, few-shot, and chain-of-thought configurations.
(4) Long Context: Ability to process and evaluate tasks involving extended input contexts (e.g., exceeding 4,000 tokens).
(5) Tool Use: Evaluating LLMs exhibiting agent-like behavior, including tool use, external API calls, and autonomous multi-step decision-making.
(6) Distributed Inference: Support for parallelized evaluation across multiple processes or compute nodes.
(7) Visual Analysis: Generation of interpretable visualizations, including bar charts, radar plots, and heatmaps, to facilitate comparative analysis.
(8) Multilingual (Indic): Native support for evaluation on benchmarks in Indic languages.
(9) Production Optimization: Implementation of runtime optimizations such as batching, caching, and prompt reuse to enhance evaluation efficiency.
(10) Interactive CLI: Availability of a command-line interface for interactive configuration of datasets, models, prompting strategies, and visualizations.
(11) Quantization: Compatibility with quantized model formats (e.g., 8-bit, 4-bit weights) to minimize memory requirements.
To contextualize the contributions of Eka-Eval, we benchmark it against established alternatives. Table 1 presents a comparative analysis of Eka-Eval and seven widely used LLM evaluation frameworks across the eleven capabilities outlined in this work. Eka-Eval demonstrates robust support across all key capabilities.
4 Eka-Eval
4.1 Design and Implementation
Eka-Eval is architected as a modular, extensible evaluation framework that balances comprehensive benchmark coverage with practical usability. The system is designed around three core principles: modularity for easy extension and customization; accessibility for low-barrier adoption across diverse research environments; and comprehensiveness supports a wide range of capabilities, including underserved areas such as long-context reasoning and tool use.
4.2 System Architecture
The framework follows a layered architecture consisting of four primary components:
4.2.1 Evaluation Engine
This component orchestrates all evaluation workflows:
Task Scheduler: Manages task scheduling, prompt formatting, and result aggregation across distributed inference setups. The scheduler implements intelligent work distribution as demonstrated in main_orchestrator() - dynamically assigning evaluation tasks to available workers based on resource constraints and model requirements.
Batch Optimizer: Implements intelligent batching strategies and supports various quantization schemes to optimize memory usage and inference speed. As seen in the PIQA evaluation implementation, the optimizer automatically adjusts generation_batch_size parameters to maximize throughput while preventing out-of-memory errors.
Distributed Coordinator: Coordinates evaluation across multiple GPUs and workers using Python’s multiprocessing library. The coordinator launches multiple worker_process instances to handle independent evaluation tasks, enabling parallel execution across different benchmarks and model configurations.
4.2.2 Benchmark Registry
Provides a unified interface for managing datasets:
Dataset Manager: The BenchmarkRegistry class handles diverse dataset formats and sources, abstracting the complexities of different evaluation protocols. The manager supports datasets from HuggingFace Hub, local files, and custom data formats through standardized interfaces.
4.2.3 Model Interface Layer
Abstracts access to different local and API-based model backends:
Local Model Loader: Initializes transformer-based checkpoints with automatic device allocation and quantization.
API Client Manager: Manages proprietary endpoints through dedicated clients (OpenAIClient, GeminiClient, ClaudeClient) that extend BaseAPIClient, providing unified request handling with rate limiting and authentication.
Interactive Selection Interface: Implements get_model_selection_interface() for dynamic model discovery and selection, supporting local model paths and API configurations.
Resource Manager: Ensures efficient memory management through explicit cleanup functions, preventing resource leaks during repeated evaluation runs.
Framework Setup & Config Navigation Result Export Indic Support Extensibility
[…]
4.2.4 Results Processing System
Handles comprehensive output management through three secondary components:
Metrics Calculator: Computes evaluation metrics using HuggingFace’s evaluate library (e.g., accuracy, BLEU, F1-score, exact match, Pass@1). It also implements robust error handling for edge cases and missing data; for example, when a model returns “The answer is probably B”, regex-based extraction retrieves the label; if that fails, a default score is assigned. Code completions are sandboxed with timeout control to ensure safe execution and error logging.
Visualisations analytics: Provides comparative analysis across multiple models and benchmark configurations by generating visualizations such as bar charts, heatmaps and radar plots (including support for cross-model performance comparisons).
Export Manager: Handles result export in multiple formats including JSON and CSV. The manager maintains evaluation metadata including model parameters, benchmark versions, execution timestamps, and system configurations.
4.3 Comprehensive Benchmark Coverage
Eka-Eval covers nine major evaluation categories with comprehensive benchmark support across 35+ benchmarks (See Appendix A). These categories include:
  1. 1.
    Code Generation and Programming: Programming abilities are assessed using HumanEval (Chen et al., 2021), MBPP  (Austin et al., 2021), HumanEval+ (Liu et al., 2023), EvalPlus (Liu et al., 2023), with Pass@1 accuracy metrics.
  2. 2.
    Mathematics and Logical Reasoning: Mathematical capabilities are evaluated through GSM8K (Cobbe et al., 2021) for grade school math, MATH (Hendrycks et al., 2021) for competition-level problems, and ARC-C (Clark et al., 2018) for scientific reasoning.
  3. 3.
    Reading Comprehension: Text understanding capabilities are evaluated using SQuAD Rajpurkar et al. (2018), QuAC Choi et al. (2018), and BoolQ Clark et al. (2019) with F1 scores and exact match metrics.
  4. 4.
    Commonsense Reasoning: We incorporate PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), ARC-C (Clark et al., 2018), WinoGrande (Sakaguchi et al., 2021), CommonSenseQA (Talmor et al., 2018), and OpenBookQA (Mihaylov et al., 2018) for comprehensive commonsense evaluation.
  5. 5.
    World Knowledge: Factual knowledge is tested through TriviaQA (5-shot) (Joshi et al., 2017) and NaturalQuestions (5-shot) (Kwiatkowski et al., 2019) with accuracy metrics.
  6. 6.
    Long-Context Understanding: For extended context reasoning, we include ZeroSCROLLS (Shaham et al., 2023) with ROUGE (Lin, 2004) and F1 scores, Needle-in-a-Haystack (Wang et al., 2024a) for retrieval accuracy, and InfiniteBench (Zhang et al., 2024) for task-specific long-context evaluation.
  7. 7.
    General Reasoning and Knowledge: For foundational capabilities, we include MMLU (Hendrycks et al., 2020a) and MMLU-Pro (Wang et al., 2024b) for multitask language understanding, IFEval (Zhou et al., 2023) for instruction following, BBH (3-shot) (Suzgun et al., 2022) for challenging reasoning tasks, and AGI-Eval (3-5 shot) (Zhong et al., 2023) for general intelligence assessment.
  8. 8.
    Tool Use and API Reasoning: Practical capabilities are assessed through API-Bank Li et al. (2023) for API call accuracy and ROUGE-L and API-Bench Patil et al. (2024) for API recommendation accuracy.
  9. 9.
    Multilingual and Indic Language Support: A distinguishing feature of Eka-Eval is its dedicated support for Indian languages, addressing a critical gap in existing evaluation frameworks. The system includes benchmarks covering the majority of Indian languages. For multilingual evaluation, it supports popular Indic language benchmarks including MMLU-IN (Hendrycks et al., 2020b), TriviaQA-IN (Joshi et al., 2017), MILU (Verma et al., 2024), GSM8K-IN (Cobbe et al., 2021), BoolQ-IN (Clark et al., 2019), ARC-C-IN (Clark et al., 2018), Flores-IN (Goyal et al., 2022) with BLEU (Papineni et al., 2002) and ChrF metrics (Popović, 2015), XQuAD-IN (Artetxe et al., 2019) and XorQA-IN (Asai et al., 2021) with F1 and exact match scores.
    The multilingual architecture includes language-specific prompts, culturally appropriate protocols, and specialized tokenization for Indic scripts, ensuring fair evaluation across languages while staying consistent with global benchmarks.
4.4 Extensibility and Customization
The framework is designed with extensibility as a first-class concern. New benchmarks can be integrated through a simple plugin architecture that requires minimal boilerplate code. The system supports custom evaluation metrics, prompt templates, and post-processing pipelines, enabling users to adapt the framework to their specific requirements.
Configuration is managed through a hierarchical JSON-based system that enables users to define evaluation suites ranging from quick smoke tests to comprehensive benchmark runs. It also supports parameter sweeps, facilitating systematic exploration of prompt variations, few-shot examples, and model hyperparameters.
5 Experiments
To assess the effectiveness and usability of Eka-Eval, we conducted a comprehensive evaluation combining benchmark coverage analysis and user-centered feedback across six prominent frameworks: lm-eval-harness, OpenCompass, HELM, FreeEval, indic-eval, and Eka-Eval.
5.1 Evaluation Procedure
Three human participants independently evaluated each framework by running the Sarvam-1B Sarvam (2024) model on four diverse benchmarks: WinoGrande, PIQA, ARC-C and FLORES. Participants installed each framework in a standardized environment, integrated the target model, and ran the recommended benchmark workflows. They then rated the framework on five criteria using Likert scale ranging from 1 (Very Poor) to 5 (Excellent):
EKA-EVAL Analysis: All ratings were recorded to enable direct comparison across frameworks. Detailed instructions and a standardized rating form were provided to ensure consistency and reproducibility. Participants were encouraged to consult official documentation but not to seek external technical support beyond publicly available resources.
Participants assessed each framework across five evaluation criteria:
  1. 1.
    Setup and Configuration Time: Time and effort required to install dependencies, configure models, and execute an initial benchmark run.
  2. 2.
    Ease of Navigation: Intuitiveness of navigation, benchmark selection, and configuration, including CLI clarity, documentation quality, and ease of discovering options.
  3. 3.
    Result Reporting and Export: Clarity and accessibility of evaluation outputs, with options to export results (e.g., JSON, CSV) and create visualizations such as bar charts or heatmaps.
  4. 4.
    Indic Language Support: Support for Indic and multilingual datasets such as ARC-IN and FLORES. Evaluators assessed the availability of prebuilt resources.
  5. 5.
    Extensibility: Ease of customizing the framework to add new prompt templates, models, datasets, or evaluation metrics.
[…]
6 Conclusion and Future Work
In this work, we introduced Eka-Eval, a unified and extensible framework designed to streamline the evaluation of LLMs across 35+ diverse benchmarks and 11 key capabilities, spanning nine major evaluation categories-including support for Indic languages. Through a combination of system design, implementation, and user-centered evaluation, Eka-Eval demonstrated high usability and practicality, achieving the highest participant ratings in four out of five evaluation criteria. These results highlight its effectiveness in enabling reproducible, scalable, and inclusive evaluation workflows.
We plan to expand the framework to over 100 tasks, with a focus on underrepresented Indic languages. Future work will categorize benchmarks by language diversity, task difficulty, and domains such as law, healthcare, and governance. We plan to add dynamic task calibration for context length, ambiguity, reasoning complexity, and India-specific knowledge tasks. Priorities include benchmarks for bias detection, hallucination analysis, privacy risks, domain-specific assessments, and adversarial human review. Finally, we will support phase-wise evaluation across pretraining, fine-tuning, and deployment to provide a comprehensive view of model behavior. These enhancements will be included in Eka-Eval v2.0, strengthening its role as the evaluation standard for Indic and other LLMs.
7 Ethics Statement
This research uses publicly available data without personally identifiable information. All datasets and models comply with their terms of use. The work is intended for academic research. Potential misuse or unintended amplification of biases should be carefully considered before deployment.
Appendix A Appendix
A.1 Indic benchmark configuration
As demonstrated in Figure 2, a sample configuration used to evaluate the ARC-Challenge-Indic benchmark across 11 Indic languages. It illustrates how task parameters, templates, and dataset references are modularly specified in Eka-Eval.
"ARC-Challenge-Indic": {
  "description": "Zero-shot evaluation
    across 11 Indic languages",
  "evaluation_function":
    "indic.arc_c_in.evaluate_arc_c_in",
  "task_args": {
    "dataset_name":
      "sarvamai/arc-challenge-indic",
    "target_languages": ["bn", "en",
      "gu", "hi", "kn", "ml", "mr",
      "or", "pa", "ta", "te"],
    "dataset_split": "validation",
    "num_few_shot": 0,
    "max_new_tokens": 10,
    "generation_batch_size": 8,
    "prompt_template_name_zeroshot":
      "arc_c_in_0shot",
    "prompt_template_name_fewshot":
      "arc_c_in_5shot",
    "prompt_file_benchmark_key":
      "arc_c_in",
    "prompt_file_category": "indic"
Figure 2: ARC-Challenge-Indic benchmark configuration example
A.2 CLI Demonstration
The interactive CLI of the Eka-Eval framework is shown below, which guides users through model selection and evaluation setup. Simplifying benchmarking workflows, it is accessible to both researchers and developers.
Figure 3: Available benchmarks groups of Eka-Eval framework.
As per Figure 3, users are prompted to select high-level task groups (e.g., Reading Comprehension) during CLI setup. This enables fine-grained benchmarking organization and streamlined selection.
Figure 4: Model selection in the Eka-Eval framework.
Eka-Eval supports local HuggingFace models and API-based models like Sarvam, Gemma, OpenAI, Claude, and Gemini. Users interactively select model source and configuration through CLI. (see Figure 4)
Figure 5: Subtask selection within a task group.
After selecting a task group, users choose specific benchmarks such as SQuAD, BoolQ, or QuAC for focused evaluation within that domain. (see Figure 5)
Figure 6: Consolidated evaluation results table.
The CLI displays final benchmark scores for each model in tabular format, including per-task and average scores. Results are also exported as CSV. (see Figure 6)
Figure 7: Interactive visualisation setup in Eka-Eval.
As per Figure 7) framework allows users to generate multiple types of visualizations: bar charts, heatmaps, radar plots that are based on completed evaluations.
Figure 8: Bar chart visualisation
Figure 8 shows performance breakdown across sub-tasks like BoolQ, SQuAD, and QuAC. It provides intuitive insight into strengths and weaknesses of the model.
A.2.1 Prompt Template System
A critical component of Eka-Eval is its sophisticated prompt management system, which handles diverse evaluation paradigms and languages. The framework implements a flexible template system demonstrated through PIQA benchmark prompt 9:
  "piqa_generation": {
    "template": "Choose the most appropriate
      solution (0 or 1) to achieve the goal:
[…]
    "description": "Generation-based PIQA prompt"
  "piqa_5shot_generation": {
    "template_prefix": "Choose the most
      appropriate solution (0 or 1)...",
    "few_shot_example_template":
[…]
    "template_suffix": "Question: {goal}
[…]
    "description": "Few-shot generation prompt"
  "default_few_shot_examples_piqa": [
      "goal": "To remove a stain from clothing",
      "sol1": "Apply cold water immediately...",
      "sol2": "Set the clothing on fire...",
      "answer_label": "0"
Figure 9: PIQA prompt templates supporting multiple evaluation paradigms
The Prompt template system as shown in Figure 9 supports zero-shot, few-shot, and chain-of-thought prompting strategies, ensuring consistency across evaluation modes and languages. Users can customize prompt strategies and easily configure them in the benchmark configuration file, as shown in Figure 2.
Appendix B Limitations
While Eka-Eval supports a wide range of benchmarks and model backends, it currently lacks a graphical user interface, relying instead on CLI-based workflows. The framework does not yet support vLLMs and has limited support for detailed error analysis or explainability. Additionally, reproducibility may be affected by changes in external datasets or model versions if not explicitly versioned or cached.
Appendix C Acknowledgements
The authors express their gratitude to Ritika Iyer, Nikhil Mishra, and Rajat Kumar Thakur for their contributions in evaluating the framework, reviewing the manuscript, and reporting the results. We also appreciate the valuable suggestions and feedback provided by Aamod Thakur, Prathamesh Shanbhag and Shailesh Panda.